OpenStudy (johnkim2330):

HELP WILL GIVE MEDAL What is the simplified form of the quantity x over 3 plus y over 4 all over the quantity x over 4 minus y over 3 ?

OpenStudy (johnkim2330):

\[\frac{ \frac{ X }{ 3 }-\frac{ Y }{ 4 } }{ \frac{ X }{ 4 } -\frac{ Y }{ 3 }}\]

OpenStudy (johnkim2330):

CHOICES ARE the quantity 4x plus 3y all over the quantity 3x minus 4y the quantity 4x minus 3y all over the quantity 3x plus 4y the quantity 3x plus 4y all over the quantity 4x minus 3y the quantity 3x minus 4y all over the quantity 4x plus 3y

OpenStudy (johnkim2330):

@shiraz14

OpenStudy (shiraz14):

Sorry, I'm not interested ... bye.